Grade 9Algebra

ax + Given that a, b, and c are non-zero real numbers and a + b ≠ 0, solve for x. ax + bx - c = 0

tsmoke
7 Years agoGrade 9

1 Answer

Monoxdifly
7 Years ago
ax + bx – c = 0
Add each hand by c:
ax + bx – c + c = 0 + c
ax + bx = c
Use distributive property:
(a + b)x = c
Divide both hands by (a + b):
x = c/(a + b)